OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (anonymous):

\[u\cdot v=\|u\|~\|v\|~\cos\color{red}{\theta}\] where \(u\cdot v\) is the dot product of the two vectors, and \(\|u\|\) is the magnitude of \(u\).